### v1

# SLIM: Let LLM Learn More and Forget Less with Soft LoRA and Identity Mixture

Abstract
Although many efforts have been made, it is still a challenge to balance the training budget, downstream performance, and the general capabilities of the LLMs in many applications. Training the whole model for downstream tasks is expensive, and could easily result in catastrophic forgetting. By introducing parameter-efficient fine-tuning (PEFT), the training cost could be reduced, but it still suffers from forgetting, and limits the learning on the downstream tasks. To efficiently fine-tune the LLMs with less limitation to their downstream performance while mitigating the forgetting of general capabilities, we propose a novel mixture of expert (MoE) framework based on Soft LoRA and Identity Mixture (SLIM), that allows dynamic routing between LoRA adapters and skipping connection, enables the suppression of forgetting. We adopt weight-yielding with sliding clustering for better out-of-domain distinguish to enhance the routing. We also propose to convert the mixture of low-rank adapters to the model merging formulation and introduce fast dynamic merging of LoRA adapters to keep the general capabilities of the base model. Extensive experiments demonstrate that the proposed SLIM is comparable to the state-of-the-art PEFT approaches on the downstream tasks while achieving the leading performance in mitigating catastrophic forgetting.
SLIM: Let LLM Learn More and Forget Less with Soft LoRA and Identity Mixture
Jiayi Han1,2, Liang Du3, Hongwei Du1,2, Xiangguo Zhou1,2, Yiwen Wu1,2, Weibo Zheng1,2, Donghong Han4, 1Inspur Genresoft Co. Ltd., Inspur Group Co. Ltd. 2Shandong Key Laboratory of Automated Complex Network Software Construction 3Interactive Entertainment Group, Tencent Inc. 4Northeastern University, China Correspondence: hanjiayi@inspur.com, handonghong@cse.neu.edu.cn
1 Introduction
In recent years, large language models (LLMs) have demonstrated extraordinary general capabilities, and are widely used for question-answering, code generation, mathematics, and other fields. However, the increasing utilization of LLMs in specialized domains needs to support tasks with a mixture of world and specific knowledge. Meanwhile, since the scale of the LLMs is large, fine-tuning the whole model could be expensive, and may not be affordable for many users. Furthermore, after fine-tuning on the downstream tasks, the LLMs may lose the general capability, due to catastrophic forgetting. How to reduce the training cost while mitigating catastrophic forgetting without loss of learning capacity, is an essential problem in the utilization of LLMs, especially when utilized in a scenario where both common and professional capabilities are needed.
Figure 1: The radar chart of the performance of the parameter-efficient fine-tuning methods. The blue benchmarks are evaluated after fine-tuning on other downstream datasets. Compared with the state-of-the-art parameter-efficient methods, the proposed SLIM could achieve comparable downstream performance while mitigating catastrophic forgetting.
An intuitive solution is to train only a small bunch of parameters. For example, low-rank adaptation (LoRA) Hu et al. (2022), a parameter-efficient fine-tuning (PEFT) approach, is validated to forget fewer pieces of knowledge compared with fine-tuning the full weight of the base model, with limitations of learning capacity Biderman et al. (2024). Chaudhry et al. (2020) propose to optimize the model in the orthogonal subspace with low-rank adaptation to alleviate forgetting. Singh et al. (2020) propose to insert calibrating modules to the base model for learning the downstream tasks. Liang et al. (2023) utilize learnable mask prompts to fine-tune the CLIP without loss of text-vision alignment. These approaches attempt to introduce minimal changes to the base model to keep its raw capacity. Some approaches utilize data re-play to reduce forgetting. For example, Glorioso et al. (2024) utilize a 60% data replay for continual training. Zhu et al. (2024) utilize a 30% re-play rate in the continual pre-training for the coding model. Despite the fact that data replay could effectively reduce the influence of forgetting, the large amount of training data is still a burden, and the pre-training data may not be accessible for downstream fine-tuning.
In this work, we propose a novel MoE architecture with Soft LoRA and Identity Mixture (SLIM), a mixture of LoRA adapters and identity layers. By introducing identity layers as experts, the model has the potential of routing the samples out of the downstream distribution to the identity layers, thus avoiding the influence of downstream tasks. To encourage the model to dynamically route the samples from the downstream tasks to the LoRA adapters, while assigning the samples out of the downstream distribution to the identity layers to avoid forgetting, we propose weight yielding with sliding clustering to estimate the distribution of the samples from the downstream tasks and modify the routing weight according to the consistency between the input sample and the distribution. Inspired by DARE Yu et al. (2024), we further convert the low-rank adaptation to model merging task and propose a fast merging approach that dynamically merges the mixture of LoRA adapters to the base model. We conduct experiments on various downstream datasets while validating the model performance on multiple general tasks. Extensive experiments show that the proposed SLIM could achieve SOTA PEFT performance, and significantly mitigate catastrophic forgetting, as demonstrated in Fig. 1. Our main contributions could be summarized as follows:
  1. 1.
    We propose SLIM, a novel MoE PEFT algorithm that achieves comparable performance to the SOTA PEFT methods, while effectively alleviating catastrophic forgetting.
  2. 2.
    We propose weight yielding with sliding clustering that enhances the dynamic routing between the identity layers and the LoRA adapters for better downstream performance and catastrophic forgetting suppression.
  3. 3.
    We propose dynamic merging that converts the MoE low-rank adaptation to model merging, which effectively eliminates catastrophic forgetting without any data replay.
2 Related work
2.1 Parameter-efficient fine-tuning (PEFT)
With the growth of the scale of the deep models, it becomes expensive to fine-tune all the parameters of the models. Therefore, parameter-efficient fine-tuning has attracted more attention from researchers in recent years. There are several strategies of PEFT. The most common strategies are LoRA-styled, adapter insertion, and prefix-tuning.
Prefix tuning
The prefix tuning introduces extra learnable prompts and re-parameterization to the transformer layers to introduce adaptations to the original model Li and Liang (2021). P-tuning v2 removes the re-parameterization strategy, and manages to expand it to more tasks Liu et al. (2022). Some approaches, such as p-tuning Liu et al. (2023) and prompt-tuning Lester et al. (2021), propose only assigning learnable tokens to the initial word embedding layer for training efficiency. SMoP Choi et al. (2023) mixing short soft prompts for efficient training while maintaining performance gains typically induced by longer soft prompts.
Adapter modules
Adapter modules introduce additional adaptive layers to the pre-trained model for fine-tuning. Serial Adapter Houlsby et al. (2019) adding adapter modules cascaded to the self-attention layers and the FFN layers. AdapterFusion Pfeiffer et al. (2021) further proposed to insert the adapters only after the normalization layers for the FFN layers to further boost training efficiency. Some approaches also introduce parallel adapters to the transformer layers He et al. (2021); Lei et al. (2023).
LoRA-styled approaches
LoRA-styled adapters assume that the change of the full-weight matrix could be approximated by a low-rank matrix Hu et al. (2022). AdaLoRA Zhang et al. (2024) proposes to adaptively estimate the parameters’ importance and modify the adapters’ rank. DoRA Liu et al. (2024) further decomposes the pre-trained weights to scale and direction, proposing only to update the direction of the pre-trained weights. MoRA Jiang et al. (2024b) proposes to increase the rank of the adapters with the same number of parameters. Inspired by the Mixtral 8$\times$7B model Jiang et al. (2024a), LoRAMoE Dou et al. (2024) and MixLoRA Li et al. (2024) propose to mixing the LoRA adapters with dynamic routing. LoRAMoE further introduces a contrastive loss function to mitigate knowledge forgetting during downstream training.
2.2 Task incremental learning
Task incremental learning, a subfield of continual learning, addresses the challenge of learning multiple tasks sequentially without forgetting previously acquired knowledge. Many explorations have been made to mitigate knowledge forgetting.
Replay-Based methods
Replay-based methods, such as Experience Replay (ER) and Generative Replay, mitigate forgetting by reintroducing samples from previous tasks during training on new tasks. For example, Rolnick et al. (2019) leverages behavior cloning and off-policy learning from the replay for training stability. Chaudhry et al. (2019) maintain a memory bank during training, and sample from the bank when training the consequent batches. Many approaches further modify the hyperparameters and training schema (e.g. learning rate, re-warm up) along with the data re-play during the continual pre-training Zhu et al. (2024); Glorioso et al. (2024); Ibrahim et al. (2024).
Regularization-based methods
Some approaches adopt extra regularization terms to the adapted weights to mitigate forgetting. For example, Chaudhry et al. (2020) proposes to fine-tune the base model on the orthogonal sub-space of its parameters to avoid the influence of the tasks trained previously. LoRAMoE Dou et al. (2024) proposes to leverage contrastive loss during fine-tuning to alleviate the loss of world knowledge. Wang et al. (2021) introduce Adam-NSCL that forces the update of the network parameters to lie in the null space of the previous tasks for capability reservation.
3 Method
Figure 2: The overall framework of the proposed approach. “Iden” represents the identity layer. The input series are first sent to the router to predict the routing weight of each identity layer and LoRA adapters. We also calculate the distance from the input to the clusters, and the routing weight is modified with the distance by the weight yielder. Then we select top-$K$ experts according to the yielded weight. The final output is achieved by the dynamic merging that replaces the normal weighted sum of the MoE methods.
3.1 Overview of LoRA-styled adapters
For an MLP layer with weight matrix $W\in\mathbb{R}^{d_{1},d_{2}}$, a LoRA adapter utilizes two low-rank learnable matrices $B\in\mathbb{R}^{d_{1},r}$ and $A\in\mathbb{R}^{r,d_{2}}$ to modify the output as:
$\begin{array}[]{l}y=Wx+b+\text{LoRA}(x)=Wx+b+BAx,\end{array}$ (1)
in which $r<<\text{min}(d_{1},d_{2})$, and $b$ represents the bias of the MLP layer.
Although LoRA is efficient and does not directly modify the weight of the base model, it still influences the output and therefore impacts the generalization of the LLMs. Since LoRA adapters are independent modules of the base model, a direct solution for recovering the generalization is to discard the LoRA adapters from the base model when not needed. We formulate this alternative implementation of LoRA adapters as in Eq. 2:
$\begin{array}[]{l}y=(W+\hat{R}BA+(1-\hat{R})\textbf{0})x+b,\end{array}$ (2)
in which $\hat{R}:\mathbb{R}^{C}\rightarrow\{0,1\}$ is an expert (e.g., human) routing operation that determines if the LoRA adapter should be discarded. Since it is not feasible to let users decide whether to utilize the LoRA adapter for each individual instruction, we propose a dynamic approach as illustrated in the following.
3.2 Mixture of experts as soft $\hat{R}$
Similar to the MoE architecture of the Mixtral models, the mixture of LoRA adapters also introduces a routing layer to assign the input tokens to multiple LoRA adapters. Formally, given input $x\in\mathbb{R}^{B,N,C_{in}}$ in which $B,N,C_{in}$ represent the batch size, length of input tokens, and the number of channels, a routing layer $R$ is adopted to assign the tokens to the LoRA experts as in Eq. 3:
$\hat{w}_{i}=\left\{\begin{array}[]{l}\text{Softmax}\left(R(x)\right)_{i},\mathcal{R}(R(x)_{i})\leq 2\\ 0,otherwise\end{array}\right.$ (3)
in which $\mathcal{R}(R(x)_{i})$ represent the descending order of $R(x)_{i}$ in $R(x)$. Then select the top-$K$ LoRA experts to process the input as in Eq. 4:
$\left\{\begin{array}[]{l}y=Wx+\frac{1}{Z}\sum\limits_{i}\hat{w}_{i}\text{LoRA}_{i}(x)+b\\ Z=\sum_{i}\textbf{1}(\mathcal{R}(\hat{w}_{i})\leq 2)w_{i}\end{array}.\right.$ (4)
To automatically assign the tokens to the adapters, we replace the expert routing operation $\hat{R}$ with a learnable routing layer conditioned on the input as:
$\begin{array}[]{l}y=(W+\hat{R}(x)BA+(1-\hat{R}(x))\textbf{0})x+b.\end{array}$ (5)
To be more general, when multiple experts are introduced, Eq. 5 could be formulated as follows:
$\begin{array}[]{l}y=Wx+b+\sum\limits_{i}\hat{R}(x)_{i}f_{i}(x),\end{array}$ (6)
in which $f_{i}$ could either be a LoRA adapter or an identity layer that satisfies $f(x)=x$, and the routing weight $\hat{R}(x)$ satisfies $\sum_{i}\hat{R}(x)_{i}=1$. Note that we do not utilize the 0 matrix in our method. By setting $\hat{R}(x)_{i}=\hat{w}_{i}/Z$, Eq. 6 could be transformed to the same formulation with Eq. 4.
3.3 Weight yielding with sliding clustering
To identify the samples occluded from the target domain, we suppose the distribution of the input samples is a mixture of Gaussian distribution. We calculate the distance from the input sample to the nearest cluster. We suppose that if the distance is large, the sample is rarely probable to be sampled from the target domain, therefore it should be processed by the identity expert. During training, we assign each input sample to a cluster and update its center. We also estimate the variance of each channel of the cluster. We do not update the clusters during inference.
Formally, we randomly initialize $N$ cluster centers $\mathcal{C}=\{c_{i}\in\mathbb{R}^{C}\}$. For each input sample $x\in\mathbb{R}^{N,C}$, we assign the sample to the nearest cluster as follows:
$\begin{array}[]{l}\text{idx}|_{x}=\mathop{\text{argmin}}\limits_{i}\left(||c_{i}-\frac{1}{N}\sum\limits_{k}x_{k,:}||_{2}\right).\end{array}$ (7)
We then yield the routing weights as follows:
$w_{i}=\\ \left\{\begin{array}[]{l}\tilde{R}(x)_{i},\text{TYPE}(f_{i})=\text{LoRA}\\ \tilde{R}(x)_{i}+d-1,otherwise\\ \end{array},\right.$ (8)
in which $\text{TYPE}(\cdot)$ return the type of the layer, $\tilde{R}:\mathbb{R}^{C}\rightarrow\mathbb{R}$ calculate the routing logits of the input, and the distance $d$ could be calculated as follows:
$\begin{array}[]{l}d=\left(\text{max}(3-\sum\frac{|x-c_{\text{idx}}|}{\sigma_{\text{idx}|_{x}}C},1e-4)\right)^{-1}.\end{array}$ (9)
The routing weight could finally be formulated as follows:
$\hat{w}_{i}=\left\{\begin{array}[]{l}\text{Softmax}(w_{i}),\mathcal{R}(w_{i})\leq K\\ 0,otherwise\end{array},\right.$ (10)
in which $K$ represents the number of experts to be activated.
3.4 Mixing LoRA adapters with dynamic merging
DARE Yu et al. (2024) proposes to merge the LLMs with the same pre-trained parameters to obtain the capacity of the merged LLMs. DARE supposes the weights of the fine-tuned models could be formulated as $\theta_{\text{SFT}}=\theta_{\text{PRE}}+\tau$, in which $\theta_{\text{SFT}},\theta_{\text{PRE}}$ represent the weight of the fine-tuned model and the pre-trained model. $\tau$ is named as “task vector”. As depicted in DARE, LLM have the following natures:
  1. 1.
    LLM is robust to slight parameter changes.
  2. 2.
    By merging the task vectors of fine-tuned LLMs, the LLM could obtain multiple downstream capabilities.
Inspired by the findings of DARE, we propose to convert the mixing of multiple LoRA adapters to dynamic model merging to further recover the generalization of the LLM. DARE randomly sets the elements of $\tau$ to 00 to formulate $\theta_{\text{DARE}}$. The merging of the model could be formulated as follows:
$\theta_{M}&=\theta_{\text{PRE}}+\lambda\sum\limits_{k}(\theta_{\text{DARE}}^{t_{k}}-\theta_{\text{PRE}})\\ &=\theta_{\text{PRE}}+\frac{\lambda}{1-p}\sum\limits_{k}M_{t_{k}}(\theta^{t_{k}}_{\text{SFT}}-\theta_{\text{PRE}})\end{aligned},$ (11)
in which $\otimes$ represents the Hadamard product, and $p\in(0,1)$ is the masking rate. We formulate the merging of the general instruction tuning model and the downstream model trained by LoRA as:
$\theta_{M}&=\theta_{\text{PRE}}\\ &+\frac{\lambda}{1-p}M_{L}\otimes(\theta_{\text{Ins}}-\theta_{\text{PRE}})\\ &+\frac{\lambda}{1-p}M_{A}\otimes(\theta_{\text{Ins}}+BA-\theta_{\text{PRE}})\\ \end{aligned},$ (12)
in which $\theta_{M},\theta_{\text{PRE}},\theta_{\text{Ins}}$ represent the weight of the merged model, the pre-trained LLM, and the instruction tuning model, $M_{L}$ and $M_{A}$ are binary masks. To simplify the merging procedure, we need to eliminate the $\theta_{\text{PRE}}$ from the equation to avoid access to the pre-training model during inference. Suppose $M_{L}$ and $M_{A}$ satisfy $XOR(M_{L},M_{A})=\textbf{1}$. Setting $\lambda=1-p$, Eq. 12 could be formulated as follows:
$\theta_{M}&=\theta_{\text{PRE}}+\frac{\lambda}{1-p}M_{L}\otimes(\theta_{\text{Ins}}-\theta_{\text{PRE}})\\ &+\frac{\lambda}{1-p}M_{A}\otimes(\theta_{\text{Ins}}+BA-\theta_{\text{PRE}})\\ &=\theta_{\text{Ins}}+M_{A}\otimes(BA)\end{aligned}.$ (13)
Figure 3: The FAST implementation of dynamic merging. By masking the sub-matrices of the low-rank adapter, we approximate the random sampling without introducing extra computational cost.
Then formulate the mixture of LoRA experts as:
$\begin{array}[]{l}\theta_{M}=\theta_{Ins}+\sum_{i}\hat{w}_{i}M_{A}^{i}\otimes(B_{i}A_{i})\end{array}.$ (14)
It is worth noting that we ignore the identity layers for simplicity. Although Eq. 14 could mimic the DARE approach to the mixture of LoRA experts, calculation of the full $BA$ for all LoRA adapters is required, resulting in multiple $d_{1}\times d_{2}$ matrices, which is not computationally efficient. To address this issue, we propose a fast implementation to approximate the random masking.
Formally, denote the row set of the masked elements as $\mathcal{S}_{\text{row}}=\{i|\exists j\in[0,d_{2}],M_{i,j}=1\}$ and the set of the column set $\mathcal{S}_{\text{col}}=\{j|i\in[0,d_{1}],M_{i,j}=1\}$, we could approximate the masking as follows:
$\left\{\begin{array}[]{l}B^{\prime}=\text{MASK}(B,\mathcal{S}_{\text{row}})\\ A^{\prime}=\text{MASK}(A^{T},\mathcal{S}_{\text{col}})^{T}\\ M_{A}\otimes(BA)\approx B^{\prime}A^{\prime}\end{array},\right.$ (15)
in which $\text{MASK}(a,b)$ is a function that masks the rows of $a$ according to the index set $b$, which could be formulated as follows:
$\text{MASK}(a,b)_{i,:}\left\{\begin{array}[]{l}\textbf{0},i\in b\\ a_{i,:},otherwise\end{array}.\right.$ (16)
Compared with masking with Hadamard product, this approach will set the whole column (row) to 0, which dramatically increases the masking ratio. To address this, we randomly sample a subset of $\mathcal{S}_{\text{row}}$ and $\mathcal{S}_{\text{col}}$ to fit the original masking ratio. It is worth noting that during inference, we directly masking the rows and columns, instead of the approximation approach mentioned above.
We also introduce $L_{1}$ regularization to the adapters. Denote $l$ and $k$ as the index of layers and LoRA adapters, it could be formulated as:
$\begin{array}[]{l}\mathcal{L}_{L_{1}}=\sum_{l}\sum_{k}\left(\frac{B_{k}^{l}}{r\times C_{\text{out}}}+\frac{A_{k}^{l}}{r\times C_{\text{in}}}\right)\end{array}.$ (17)
3.5 Summarising the proposed SLIM
To learn the downstream tasks while mitigating catastrophic forgetting, the proposed SLIM modifies the LoRA architecture on the following sides:
  1. 1.
    Inspired by Mixtral Jiang et al. (2024a), we propose to utilize mixture of experts architecture as soft $\hat{R}$ in Eq. 2. To mimic the discarding operation, we utilize identity layers along with the LoRA adapters as experts.
  2. 2.
    We introduce weight yielding with sliding clustering to correct the routing weight and encourage the model to predict downstream tasks with the trained adapters, and the general tasks without them.
  3. 3.
    We also propose a dynamic merging approach that converts the MoE to the model merging formulation to fuse the general capacity of the base model and the downstream capability of the fine-tuned adapters. The overall architecture is depicted in Fig. 2.
For a given input, we first assign it to its nearest cluster center and calculate the distance in between. During training, the centers and the distributions of the clusters are updated according to the assigned samples. Then the router predicts the routing weight of the input. The routing weight is then yielded according to the distance, as depicted in Eq. 9, and then normalized to the final routing weight as in Eq. 10. We activate the mixed experts including identity layers and LoRA adapters according to the routing weight, and utilizing the dynamic merging to obtain the final output. The whole process could be summarized as
$\begin{array}[]{l}y=Wx+b+\sum_{i}\hat{w}_{i}f_{i}(x),\\ f_{i}(x)=\left\{\begin{array}[]{c}x,i\leq K\\ B^{\prime}_{i}A^{\prime}_{i}x,otherwise\end{array},\right.\end{array}$ (18)
in which all symbols have the same meaning as in the equations mentioned earlier.
4 Experiments
4.1 Implementation details
We evaluate the proposed SLIM on the Openchat-8B model Wang et al. (2024), an extension of Llama3-8B-Instruct. We conduct two different training settings: single dataset setting (SDS) and multi-dataset setting (MDS). For SDS, we train the model with a single downstream task. For MDS, we mix the datasets for model training. MDS include OBQA Mihaylov et al. (2018), SIQA Sap et al. (2019) and BOOLQ Clark et al. (2019) datasets. For SDS, we involve CSQA Talmor et al. (2019), Hellaswag Zellers et al. (2019), Winogrande Sakaguchi et al. (2021), ARC-e and ARC-c Clark et al. (2018) datasets. The fine-tuned models under MDS are evaluated on the general tasks including MMLU Hendrycks et al. (2021), GSM8K Cobbe et al. (2021), and PIQA Bisk et al. (2020) to measure the catastrophic forgetting of the PEFT approaches. We fine-tune the model 2 epochs for all tasks, on a single NVIDIA-A100 80G GPU, the batch size is set to 16.
4.2 Comparison with SOTA approaches
First, we compare the proposed approach with other widely used PEFT approaches. LoRAMoE, MixLoRA, MoLA are set to $rank=16,N_{\text{experts}}=8$, and SLIM holds $N_{\text{experts}}=10$ with 2 identity layers as experts, contains the same number of parameters. For a fair comparison, we set the $rank$ of LoRA and DoRA to 128. The results are shown in Tab. 1. The result demonstrates that the proposed approach achieves comparable performance on the downstream tasks, and achieves the best performance on multiple datasets. The models fine-tuned under the MDS setting are further evaluated on the general datasets, and the results are shown in Tab. 2. Compared with the other PEFT approaches, the proposed SLIM method significantly alleviates the loss of generalization capacity of the fine-tuned model, without any extra data replay.
Table 1: Comparison with SOTA approaches on downstream tasks. We BOLD the best result for each downstream task and the average performance.
Model SDS MDS AVG
CSQA HellaS WinoG ARC-c ARC-e OBQA SIQA BOOLQ
LoRA Hu et al. (2022) 78.13 87.82 78.37 81.65 85.26 68.40 72.36 67.15 77.39
DoRA Liu et al. (2024) 78.95 87.93 79.29 81.56 82.28 69.60 71.85 67.09 77.32
LoRAMoE Dou et al. (2024) 88.78 94.61 84.21 84.21 89.68 87.00 77.53 74.98 85.12
MixLoRA Li et al. (2024) 85.01 93.68 85.08 82.33 85.48 84.00 77.89 72.11 83.06
MixLoRA-Dy Li et al. (2024) 85.50 93.82 84.92 83.19 87.83 82.60 78.40 73.30 83.16
MoLA Gao et al. (2024) 85.66 93.82 82.00 82.84 87.33 82.00 78.40 72.84 83.11
SLIM (Ours) 93.28 94.87 84.13 88.22 91.83 87.00 81.57 72.14 86.63
Table 2: Comparison with SOTA approaches on general tasks with the base model fine-tuned under MDS to measure catastrophic forgetting. We BOLD the best result for each task and the average performance.
Model Generalization AVG
MMLU GSM8K PIQA
LoRA 32.73 0.00 60.99 31.24
DoRA 31.07 0.00 57.18 29.41
LoRAMoE 60.20 59.43 79.54 66.39
MixLoRA 55.41 20.47 77.36 51.08
MixLoRA-Dy 56.14 21.38 78.43 51.98
MoLA 53.15 15.54 74.70 47.79
SLIM (Ours) 65.83 76.11 84.65 75.65
4.3 Ablation study
Effectiveness of the main modules
We first validate the effectiveness of the proposed modules, and the results are shown in Tab. 3. The baseline is the MixLoRA approach. Simply removing the LoRA adapters for the attention layers has a minor effect on the MDS dataset, but reduces the forgetting by 2.29%. The insertion of identity layers reduces the forgetting by 0.96%, but resulting in a performance drop on the MDS. The weight yielding further boosts the model by 2.59% compared with the baseline model without LoRA for attention layers. The dynamic merging significantly boots the model on MMLU by 5.54%, demonstrating its effectiveness for forgetting alleviation. Surprisingly, dynamic merging also improves the model performance on the MDS datasets, which might be because the general capability could benefit some downstream tasks.
Table 3: Ablation study of main modules of the proposed approach. “Attn. w/o LoRA” means we eliminate the LoRA adapters for the attention layers, and “Iden” represents the insertion of identity layers, respectively.
Attn. w/o LoRA Iden Weight yielding Dynamic merging MDS MMLU
$\times$ $\times$ $\times$ $\times$ 78.00 55.41
$\checkmark$ $\times$ $\times$ $\times$ 77.80 57.70
$\times$ $\checkmark$ $\times$ $\times$ 76.87 56.37
$\checkmark$ $\checkmark$ $\checkmark$ $\times$ 79.10 60.64
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 80.23 65.83
Table 4: The comparison of the standard dynamic merging (STD) and its fast implementation (FAST). The fast implementation dramatically speeds up the model, while having a minor influence on the performance.
method OBQA SIQA BOOLQ MMLU Time(ms)
FAST 87.00 81.57 72.14 65.83 12.02
STD 86.60 81.21 71.83 65.97 41.11
Efficiency of dynamic merging
To validate the efficiency and effectiveness of the proposed fast merging approach, we conduct an experiment on the MDS and the MMLU dataset, and the results are demonstrated in Tab. 4. The fast implementation of the dynamic merging has a slight influence on the model performance, while dramatically reducing the time cost.
Influence of the masking rate
As the masking ratio controls the contribution of the LoRA experts, we conduct experiments to measure this influence. As depicted in Eq. 14, we sample rows from the low-rank matrices.It is worth noting that the masking rate $p$ is quadratic to the masking rate of the rows and columns $p_{rc}$. Specifically, $p_{rc}=1-\sqrt{p}$. For convenience, we validate the masking rate of rows and columns, instead of directly validating the masking rate.
The experiment results are shown in Fig. 4. If the masking ratio is too large, the LoRA experts will be eliminated from the model, resulting in a significant performance drop in the downstream tasks. On the contrary, a small masking ratio can not inhibit catastrophic forgetting, therefore resulting in a degenerated performance for MMLU. We find that utilizing a balanced masking rate $p$ (e.g., 0.5) boosts both downstream tasks while maintaining the generalization of the model, which may be because the downstream task could somewhat benefit from the general capability of the instruction tuning model.
Degradation of general capacity during fine-tuning
To validate the general ability degradation of different methods, we evaluate LoRA, MixLoRA, and SLIM in the MDS training process on the MMLU dataset. The result is shown in Fig. 5. Compared with LoRA and MixLoRA approaches, the proposed SLIM is robust to the fine-tuning procedure, while the LoRA and MixLoRA degrade along with the training process.
Figure 4: The influence of masking ratio for the rows of $B$ and $A^{T}$ of LoRA adapters. The masking ratio is a trade-off between the downstream performance and the general capability.
Figure 5: Model evaluation on the MMLU benchmark during the fine-tuning with LoRA, MixLoRA and the proposed SLIM on the MDS.
Figure 6: The validation of the proposed SLIM on MDS (downstream) and MMLU (general) with different number of LoRA experts.
Validation for the number of LoRA adapters in SLIM
We conduct experiments on the MDS to measure the influence of the number of LoRA adapters. The result is shown in Fig. 6. A larger $N_{\text{experts}}$ enhances the downstream performance and has no obvious influence on the result of MMLU.
4.4 Case study on the GSM8K dataset
We further utilize a case study to visualize the result of DoRA, LoRAMoE, and the proposed SLIM. We set the temperature $t=0$, and repetition penalty $p=0$. For DoRA, we set $p=1.5$ since it tends to generate repeated output. The results are listed as follows:
Question: In a dance class of 20 students, 20% enrolled in contemporary dance, 25% of the remaining enrolled in jazz dance, and the rest enrolled in hip-hop dance. What percentage of the entire students enrolled in hip-hop dance?
DoRA: (An) The new student arrives tomorrow on Monday because her sister was fired from here job after working into someone else’s home last Friday night. Why did Addison do this? (A) … (B) … (C) … Answer: B
LoRAMoE: … The rest of the students enrolled in hip-hop dance, so 12 students enrolled in hip-hop dance. The answer is 12
SLIM: … The rest of the students enrolled in hip-hop dance, so 16 - 4 = 12 students. The percentage of the entire students enrolled in hip-hop dance is 12/20 x 100 = 60% The answer is 60
The results show that the DoRA tends to convert the GSM8K questions to multi-choice questions. Both LoRAMoE and SLIM are capable of generating formulated output, but LoRAMoE failed to understand the question asking about “percentage”.
5 Conclusion
In this work, we propose SLIM, an MoE parameter-efficient fine-tuning approach. In addition to implementing the LoRA experts, SLIM also adopts identity layers as experts and dynamically routes between the LoRA adapters and identity layers. We also propose to yield the routing weights according to the distances in the latent space. We further propose dynamic merging that formulates PEFT to model merging, which effectively combines the downstream capabilities with the original capabilities of the pre-trained model. The proposed approach achieves comparable performance to the state-of-the-art PEFT methods, with much fewer capability loss.
6 Limitations
6.1 Performance degradation still exists
Although the proposed approach effectively mitigates catastrophic forgetting, we could still observe a slight performance drop after fine-tuning. This might be because random sampling is not the optimal solution for the merging of LLMs, and the loss function solely for downstream fine-tuning cannot effectively alleviate knowledge loss. Advanced exploration of sampling strategy and training objectives might be solutions to this issue.
6.2 Difficulty for implementation
Many implementation and acceleration frameworks have been proposed for LLMs, such as vLLM vLLM Team and ollama ollama Team . However, despite the LoRA adapters could be re-parameterized to the base model, and vLLM support implementing multiple LoRA adapters to the same base model, it is still a challenge to implement LLMs with mixed heterogeneous adapters (LoRA adapters and identity layers). We are working on the acceleration engine for the MoE of LoRA adapters for its further utilization.